# Gradient Alignment Improves Test-Time Adaptation for Medical Image Segmentation

### v1

@@ Abstract
Although recent years have witnessed significant advancements in medical image segmentation, the pervasive issue of domain shift among medical images from diverse centres hinders the effective deployment of pre-trained models. Many Test-time Adaptation (TTA) methods have been proposed to address this issue by fine-tuning pre-trained models with test data during inference. These methods, however, often suffer from less-satisfactory optimization due to suboptimal optimization direction (dictated by the gradient) and fixed step-size (predicated on the learning rate). In this paper, we propose the Gradient alignment-based Test-time adaptation (GraTa) method to improve both the gradient direction and learning rate in the optimization procedure. Unlike conventional TTA methods, which primarily optimize the pseudo gradient derived from a self-supervised objective, our method incorporates an auxiliary gradient with the pseudo one to facilitate gradient alignment. Such gradient alignment enables the model to excavate the similarities between different gradients and correct the gradient direction to approximate the empirical gradient related to the current segmentation task. Additionally, we design a dynamic learning rate based on the cosine similarity between the pseudo and auxiliary gradients, thereby empowering the adaptive fine-tuning of pre-trained models on diverse test data. Extensive experiments establish the effectiveness of the proposed gradient alignment and dynamic learning rate and substantiate the superiority of our GraTa method over other state-of-the-art TTA methods on a benchmark medical image segmentation task. The code and weights of pre-trained source models will be available.
Corresponding authors: Y. Pan and Y. Xia.
@@ Introduction
Medical image segmentation assumes a pivotal role in computer-aided diagnosis, offering precise delineation of specific anatomical structures. Over the past years, considerable research efforts (Tajbakhsh et al. 2020; Salpea, Tzouveli, and Kollias 2022; Liu et al. 2020) have been devoted to medical image segmentation utilizing deep-learning techniques, resulting in significant progress. However, domain shift, primarily caused by variations in scanners, imaging protocols, and operators (Gibson et al. 2018; Ghafoorian et al. 2017), poses challenges for these models pre-trained on the labeled dataset (source domain) to generalize across different medical centers (target domain).
@@
Test-time Adaptation (TTA) addresses domain shift by adapting the pre-trained source model to the distribution of the target domain using test data during inference (Liang, He, and Tan 2024). TTA can be employed in a source-free and online manner, allowing the model to adapt without accessing the source data. The mainstream TTA methods are based on fine-tuning the model by constructing self-supervised auxiliary tasks to guide the model in training on the test data, enhancing its performance and adaptability (Wang et al. 2021; Sun et al. 2020; Zhang, Levine, and Finn 2022; Iwasawa and Matsuo 2021; Nguyen et al. 2023; Niu et al. 2023; Yang et al. 2022; Zhang et al. 2023; Wen et al. 2024). Wang et al. (Wang et al. 2021) proposed a test-time entropy minimization scheme to decrease the entropy of model predictions by fine-tuning the affine parameters within batch normalization layers. Zhang et al. (Zhang et al. 2023) extended the vanilla entropy minimization loss to a generalized one to better utilize the information in the test data. Niu et al. (Niu et al. 2023) filtered and removed partial noise with large uncertainty in the supervised information and employed sharpness-aware minimization to optimize the parameters of the model toward a flat minimum. These methods demonstrate effective adaptation when the supervised information is reliable but may lead to unexpected performance degradation when the supervised information is unreliable. In this paper, our purpose is to improve the optimization direction and step-size by gradient alignment and dynamic learning rate, respectively. After aligning the gradients, we can obtain more reliable supervised information, thereby facilitating more effective training of the source model.
@@
In deep learning, the objective function serves as a measure to quantify the discrepancy between a model’s prediction and the expected output. Objective functions with similar gradients generally indicate similar objectives and contribute to robust training. Recently, many methods seek to align the gradients of multiple objective functions to improve training (Gao et al. 2021; Dandi, Barba, and Jaggi 2022; Joo et al. 2023; Le and Woo 2024; Zhao et al. 2023, 2021). Gao et al. (Gao et al. 2021) devised a gradient discriminator to align the gradients from the source and target domains to improve the similarity between two distributions for better adversarial domain adaptation. Dandi et al. (Dandi, Barba, and Jaggi 2022) proposed an algorithm for federated learning that induces the implicit regularization to promote the alignment of gradients across different clients. Joo et al. (Joo et al. 2023) presented to combine both L2 and cosine distance-based criteria as regularization terms, leveraging their strengths in aligning the local gradient to produce more robust interpretations. Different from these methods, we implicitly perform gradient alignment to capture the similarity between gradients of two distinct objective functions to improve the gradient direction for test-time adaptation.
@@
The learning rate is a critical hyperparameter in deep learning, as it determines the step-size to update the model. An excessively large learning rate may hinder the model’s convergence, while an excessively small learning rate may make the model trapped in local optima. Extensive research has been conducted to dynamically adjust the learning rate to enhance model training (Yang et al. 2022; Xiao and Zhang 2021; Wan et al. 2021; Xiong et al. 2022; Yu, Chen, and Cheng 1995). Yang et al. (Yang et al. 2022) presented a dynamic learning rate by constructing a memory bank and calculating the cosine similarity scores between the feature of the current test image and previous features from the memory bank. Xiao et al. (Xiao and Zhang 2021) devised the dynamic weighted learning to dynamically weights the learning losses of alignment and discriminability to avoid excessive alignment learning and excessive discriminant learning. Xiong et al. (Xiong et al. 2022) proposed a graph-network-based scheduler to learn a specific scheduling mechanism without restrictions and control the learning rate via reinforcement learning. In this paper, our focus is on enhancing the optimization direction by aligning two distinct gradients. The cosine similarity between these two gradients can be employed as a natural metric to measure the degree of alignment, allowing our GraTa to adjust the learning rate dynamically.
@@
DIGA (Wang et al. 2023b) $76.59$ $76.81$ $74.12$ $60.70$ $70.63$ $71.77$
MedBN (Park et al. 2024) $75.36$ $77.02$ $74.37$ $60.15$ $70.65$ $71.51$
@@
DomainAdaptor (Zhang et al. 2023) $75.94$ $77.33$ $74.45$ $61.36$ $71.92$ $72.20$
SAR (Niu et al. 2023) $75.26$ $78.07$ $75.07$ $61.13$ 73.69 $72.65$
DeTTA (Wen et al. 2024) $75.15$ $78.17$ $75.27$ $61.26$ $72.10$ $72.39$
GraTa (Ours) 76.61 78.81 76.52 66.84 $73.05$ 74.36
@@
$\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i})$ $\beta$ 78.35 $78.50$ $69.24$ $52.13$ $68.94$ $69.43$
$\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\beta$ $77.97$ $78.77$ $73.90$ $56.53$ $70.81$ $71.60$
$\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\eta$ $76.61$ 78.81 76.52 66.84 73.05 74.36
@@ Results
$\mathcal{L}_{ent}(\theta-\nabla\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\eta$ $74.69$ $78.46$ $75.60$ $56.75$ $71.79$ $71.46$
$\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\eta$ 76.61 78.81 76.52 66.84 73.05 74.36
Comparing to other TTA methods. The results of our GraTa, the “No Adapt” baseline, and eight competing TTA methods are detailed in Table 1. The BN-based methods exhibit stable but suboptimal enhancements across all domains due to the absence of updating the model parameters. TENT exhibits limited performance, while other variants (i.e., DLTTA, DomainAdaptor, and SAR) enhance the performance significantly. It can be observed that our GraTa achieves the best performance across most scenarios and improves the baseline by $5.89\%$, which demonstrates the effectiveness and superiority of our GraTa. Specifically, our GraTa boosts the performance in Domain D by $12.77\%$, while other methods exhibit marginal improvement, which means our approach extends the upper limit of TTA in challenging scenarios.
@@
Linear: $\frac{1}{2}(x+1)$ $77.15$ 79.11 $76.07$ $63.79$ $72.60$ $73.74$
Sigmoid: $\frac{1}{1+e^{-x}}$ $77.31$ $79.04$ $76.01$ $64.69$ $71.84$ $73.78$
ReLU: $max(0,x)$ $75.92$ $78.29$ $75.16$ $62.14$ $72.95$ $72.89$
Softplus: $\ln{(1+e^{x})}$ 77.60 $78.81$ $75.44$ $63.04$ $71.49$ $73.28$
Ours: $\frac{1}{4}(x+1)^{2}$ $76.61$ $78.81$ 76.52 66.84 73.05 74.36
Ablation study. To evaluate the contributions of our optimization objective and dynamic learning rate, we conducted a series of ablation experiments, as shown in Table 2. It shows that (1) utilizing $\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i})$ only obtains limited performance similar to TENT; (2) introducing $\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ and $\eta$ enhances the performance by $2.17\%$ and $2.76\%$ respectively; (3) the best performance is achieved when both are utilized jointly (i.e., our GraTa).
@@
Figure 4: Cosine similarity between the pseudo gradient and empirical gradient w/ or w/o our proposed gradient alignment.
Alignment between the empirical and pseudo gradients. To demonstrate that our GraTa can align the empirical gradient and pseudo gradient more closely, we repeated the experiments on all scenarios and visualized the cosine similarity between these two gradients, where the empirical gradient is produced by the cross-entropy loss in a supervised manner, as shown in Figure 4. We reported the average results using each domain as the source domain and the remaining ones as target domains. The greater the cosine similarity, the closer the pseudo gradient aligns with the empirical gradient. The results demonstrate the effectiveness of our GraTa in improving gradient direction and also indicate that even small alignment improvements can lead to significant performance gains.
@@
$\mathcal{L}_{ent}$ $\mathcal{L}_{con}$ $76.61$ $78.81$ $76.52$ $66.84$ $73.05$ 74.36
$\mathcal{L}_{rec}$ 76.76 $78.94$ $74.74$ $66.43$ $73.11$ $74.00$
$\mathcal{L}_{rot}$ $76.58$ 79.02 $74.39$ 67.90 $73.22$ $74.22$
$\mathcal{L}_{sup}$ $76.68$ $78.99$ $74.32$ $67.82$ $73.36$ $74.23$
$\mathcal{L}_{den}$ $76.71$ $78.94$ $75.11$ $65.18$ $73.26$ $73.84$
$\mathcal{L}_{con}$ $\mathcal{L}_{ent}$ $74.69$ $78.46$ $75.60$ $56.75$ $71.79$ $71.46$
$\mathcal{L}_{rec}$ $75.27$ $78.54$ $76.36$ $60.29$ $73.64$ $72.82$
$\mathcal{L}_{rot}$ $75.19$ $78.54$ $76.34$ $56.29$ $73.84$ $72.04$
$\mathcal{L}_{sup}$ $75.20$ $78.53$ $76.53$ $57.41$ 73.88 $72.31$
$\mathcal{L}_{den}$ $75.37$ $78.52$ 76.67 $59.38$ $73.80$ $72.75$
$\mathcal{L}_{con}$ $\mathcal{L}_{rec}$ $75.18$ $77.85$ $75.13$ $62.33$ $71.93$ $72.48$
$\mathcal{L}_{ent}$ $74.47$ $77.74$ $74.71$ $59.64$ $71.77$ $71.66$
$\mathcal{L}_{rot}$ $75.74$ $78.04$ $75.12$ $60.13$ $71.87$ $72.18$
$\mathcal{L}_{sup}$ $74.58$ $77.51$ $74.91$ $62.99$ $71.97$ $72.39$
$\mathcal{L}_{den}$ $75.44$ $77.82$ $75.66$ $60.12$ $72.00$ $72.21$
$\mathcal{L}_{con}$ $\mathcal{L}_{rot}$ $75.31$ $78.01$ $75.02$ $61.66$ $71.91$ $72.38$
$\mathcal{L}_{ent}$ $75.37$ $78.17$ $75.07$ $62.14$ $72.00$ $72.55$
$\mathcal{L}_{rec}$ $75.19$ $78.10$ $74.39$ $61.57$ $72.03$ $72.25$
$\mathcal{L}_{sup}$ $75.31$ $78.08$ $75.95$ $61.24$ $71.94$ $72.50$
$\mathcal{L}_{den}$ $75.27$ $78.21$ $74.84$ $61.56$ $71.80$ $72.34$
$\mathcal{L}_{con}$ $\mathcal{L}_{sup}$ $75.36$ $78.17$ $75.27$ $61.49$ $71.87$ $72.43$
$\mathcal{L}_{ent}$ $75.51$ $77.86$ $74.67$ $59.23$ $71.66$ $71.79$
$\mathcal{L}_{rec}$ $75.26$ $78.09$ $74.53$ $59.78$ $71.77$ $71.89$
$\mathcal{L}_{rot}$ $75.17$ $77.87$ $74.92$ $58.48$ $71.92$ $71.67$
$\mathcal{L}_{den}$ $74.94$ $77.91$ $74.83$ $60.65$ $71.96$ $72.06$
$\mathcal{L}_{con}$ $\mathcal{L}_{den}$ $75.34$ $78.02$ $74.47$ $61.74$ $72.00$ $72.32$
$\mathcal{L}_{ent}$ $75.20$ $77.44$ $74.77$ $59.22$ $71.94$ $71.71$
$\mathcal{L}_{rec}$ $75.40$ $78.09$ $74.51$ $60.71$ $71.80$ $72.10$
$\mathcal{L}_{rot}$ $75.32$ $77.71$ $76.49$ $55.30$ $71.82$ $71.33$
$\mathcal{L}_{sup}$ $75.48$ $77.87$ $75.08$ $59.82$ $71.78$ $72.00$
@@
More combinations of objective functions. For further discussion, we introduced four additional loss functions, i.e., the reconstruction loss, the rotation prediction loss, the super-resolution loss, and the denoising loss. We utilized the cross-entropy loss to calculate the rotation prediction loss and the mean-square-error loss to calculate others. To construct the super-resolution loss, we downsampled the input image by a factor of four. For the denoising loss, we added the Gaussian noise sampled from a standard normal distribution to the image. We evaluated all possible combinations of these loss functions to determine the optimal one and displayed the results in Table 5. It reveals that (1) using $\mathcal{L}_{con}$ as the pseudo objective function achieves superior performance over others; (2) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the pseudo objective function results in similar but suboptimal performance since they all require an additional branch to construct the loss, making it challenging provide appropriate gradients during inference; (3) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the auxiliary objective function can enhance the performance of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$, demonstrating the effectiveness of our GraTa.

### v3

@@ Abstract
Although recent years have witnessed significant advancements in medical image segmentation, the pervasive issue of domain shift among medical images from diverse centres hinders the effective deployment of pre-trained models. Many Test-time Adaptation (TTA) methods have been proposed to address this issue by fine-tuning pre-trained models with test data during inference. These methods, however, often suffer from less-satisfactory optimization due to suboptimal optimization direction (dictated by the gradient) and fixed step-size (predicated on the learning rate). In this paper, we propose the Gradient alignment-based Test-time adaptation (GraTa) method to improve both the gradient direction and learning rate in the optimization procedure. Unlike conventional TTA methods, which primarily optimize the pseudo gradient derived from a self-supervised objective, our method incorporates an auxiliary gradient with the pseudo one to facilitate gradient alignment. Such gradient alignment enables the model to excavate the similarities between different gradients and correct the gradient direction to approximate the empirical gradient related to the current segmentation task. Additionally, we design a dynamic learning rate based on the cosine similarity between the pseudo and auxiliary gradients, thereby empowering the adaptive fine-tuning of pre-trained models on diverse test data. Extensive experiments establish the effectiveness of the proposed gradient alignment and dynamic learning rate and substantiate the superiority of our GraTa method over other state-of-the-art TTA methods on a benchmark medical image segmentation task. The code and weights of pre-trained source models are available at https://github.com/Chen-Ziyang/GraTa.
footnotetext: Corresponding author.
@@ Introduction
Medical image segmentation assumes a pivotal role in computer-aided diagnosis, offering precise delineation of specific anatomical structures. Over the past years, considerable research efforts (Tajbakhsh et al. 2020; Salpea, Tzouveli, and Kollias 2022; Liu et al. 2020; Wang et al. 2022; Tang et al. 2023; Xu et al. 2024) have been devoted to medical image segmentation utilizing deep-learning techniques, resulting in significant progress. However, domain shift, primarily caused by variations in scanners, imaging protocols, and operators (Gibson et al. 2018; Ghafoorian et al. 2017), poses challenges for these models pre-trained on the labeled dataset (source domain) to generalize across different medical centers (target domain).
@@
Test-time Adaptation (TTA) addresses domain shift by adapting the pre-trained source model to the distribution of the target domain using test data during inference (Liang, He, and Tan 2024). TTA can be employed in a source-free and online manner, allowing the model to adapt without accessing the source data. The mainstream TTA methods are based on fine-tuning the model by constructing self-supervised auxiliary tasks to guide the model in training on the test data, enhancing its performance and adaptability (Wang et al. 2021; Sun et al. 2020; Zhang, Levine, and Finn 2022; Iwasawa and Matsuo 2021; Nguyen et al. 2023; Niu et al. 2023; Yang et al. 2022; Zhang et al. 2023; Wen et al. 2024). (Wang et al. 2021) proposed a test-time entropy minimization scheme to decrease the entropy of model predictions by fine-tuning the affine parameters within batch normalization layers. (Zhang et al. 2023) extended the vanilla entropy minimization loss to a generalized one to better utilize the information in the test data. (Niu et al. 2023) filtered and removed partial noise with large uncertainty in the supervised information and employed sharpness-aware minimization to optimize the parameters of the model toward a flat minimum. These methods demonstrate effective adaptation when the supervised information is reliable but may lead to unexpected performance degradation when the supervised information is unreliable. In this paper, our purpose is to improve the optimization direction and step-size by gradient alignment and dynamic learning rate, respectively. After aligning the gradients, we can obtain more reliable supervised information, thereby facilitating more effective training of the source model.
@@
In deep learning, the objective function serves as a measure to quantify the discrepancy between a model’s prediction and the expected output. Objective functions with similar gradients generally indicate similar objectives and contribute to robust training, which is particularly important in multi-task learning. A typical approach in multi-task learning is PCGrad (Yu et al. 2020), which is developed to mitigate gradient conflicts between task gradients via replacing the gradient by its projection onto the normal plane of another gradient. In contrast, our optimization objective does not involve resolving conflicts between multiple gradients, as it does not directly optimize multiple gradients simultaneously. Instead, the goal of our GraTa is to improve the direction of pseudo gradient, thereby enhancing the model’s fine-tuning process. Recently, many methods seek to align the gradients of multiple objective functions to improve training (Gao et al. 2021; Dandi, Barba, and Jaggi 2022; Joo et al. 2023; Le and Woo 2024; Zhao et al. 2023, 2021). (Gao et al. 2021) devised a gradient discriminator to align the gradients from the source and target domains to improve the similarity between two distributions for better adversarial domain adaptation. (Dandi, Barba, and Jaggi 2022) proposed an algorithm for federated learning that induces the implicit regularization to promote the alignment of gradients across different clients. (Joo et al. 2023) presented to combine both L2 and cosine distance-based criteria as regularization terms, leveraging their strengths in aligning the local gradient to produce more robust interpretations. Different from these methods, we implicitly perform gradient alignment to capture the similarity between gradients of two distinct objective functions to improve the gradient direction for test-time adaptation.
@@
The learning rate is a critical hyperparameter in deep learning, as it determines the step-size to update the model. An excessively large learning rate may hinder the model’s convergence, while an excessively small learning rate may make the model trapped in local optima. Extensive research has been conducted to dynamically adjust the learning rate to enhance model training (Yang et al. 2022; Xiao and Zhang 2021; Wan et al. 2021; Xiong et al. 2022; Yu, Chen, and Cheng 1995). (Yang et al. 2022) presented a dynamic learning rate by constructing a memory bank and calculating the cosine similarity scores between the feature of the current test image and previous features from the memory bank. (Xiao and Zhang 2021) devised the dynamic weighted learning to dynamically weights the learning losses of alignment and discriminability to avoid excessive alignment learning and excessive discriminant learning. (Xiong et al. 2022) proposed a graph-network-based scheduler to learn a specific scheduling mechanism without restrictions and control the learning rate via reinforcement learning. In this paper, our focus is on enhancing the optimization direction by aligning two distinct gradients. The cosine similarity between these two gradients can be employed as a natural metric to measure the degree of alignment, allowing our GraTa to adjust the learning rate dynamically.
@@
DIGA (Wang et al. 2023b) $76.57$ $77.10$ $73.64$ $62.04$ $71.54$ $72.18$
MedBN (Park et al. 2024) $75.14$ $76.77$ $74.46$ $59.91$ $72.00$ $71.65$
@@
DomainAdaptor (Zhang et al. 2023) $75.88$ $77.43$ $75.66$ $58.64$ $72.38$ $72.00$
SAR (Niu et al. 2023) $75.26$ $78.07$ $75.07$ $61.24$ 73.69 $72.67$
DeTTA (Wen et al. 2024) $75.15$ $78.17$ $75.27$ $61.26$ $73.61$ $72.69$
GraTa (Ours) 76.61 78.81 76.52 66.84 $72.94$ 74.34
@@
$\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i})$ $\beta$ 78.35 $78.50$ $69.24$ $52.13$ $68.74$ $69.39$
$\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\beta$ $77.97$ $78.77$ $73.90$ $56.53$ $70.46$ $71.53$
$\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\eta$ $76.61$ 78.81 76.52 66.84 72.94 74.34
@@ Results
$\mathcal{L}_{ent}(\theta-\nabla\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\eta$ $74.69$ $78.46$ $75.60$ $56.75$ $71.86$ $71.47$
$\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ $\eta$ 76.61 78.81 76.52 66.84 72.94 74.34
Comparing to other TTA methods. The results of our GraTa, the “No Adapt” baseline, and eight competing TTA methods are detailed in Table 1. The BN-based methods exhibit stable but suboptimal enhancements across all domains due to the absence of updating the model parameters. TENT exhibits limited performance, while other variants (i.e., DLTTA, DomainAdaptor, and SAR) enhance the performance significantly. It can be observed that our GraTa achieves the best performance across most scenarios and improves the baseline by $5.87\%$, which demonstrates the effectiveness and superiority of our GraTa. Specifically, our GraTa boosts the performance in Domain D by $12.77\%$, while other methods exhibit marginal improvement, which means our approach extends the upper limit of TTA in challenging scenarios.
@@
Linear: $\frac{1}{2}(x+1)$ $77.15$ 79.11 $76.07$ $63.79$ $72.42$ $73.71$
Sigmoid: $\frac{1}{1+e^{-x}}$ $77.31$ $79.04$ $76.01$ $64.69$ $72.10$ $73.83$
ReLU: $max(0,x)$ $75.92$ $78.29$ $75.16$ $62.14$ 73.07 $72.92$
Softplus: $\ln{(1+e^{x})}$ 77.60 $78.81$ $75.44$ $63.04$ $71.68$ $73.32$
Ours: $\frac{1}{4}(x+1)^{2}$ $76.61$ $78.81$ 76.52 66.84 $72.94$ 74.34
Ablation study. To evaluate the contributions of our optimization objective and dynamic learning rate, we conducted a series of ablation experiments, as shown in Table 2. It shows that (1) utilizing $\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i})$ only obtains limited performance similar to TENT; (2) introducing $\mathcal{L}_{con}(\theta-\nabla\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$ and $\eta$ enhances the performance by $2.14\%$ and $2.81\%$ respectively; (3) the best performance is achieved when both are utilized jointly (i.e., our GraTa).
@@
Figure 4: Cosine similarity between the pseudo gradient and empirical gradient w/ or w/o our proposed gradient alignment. The objectives of “TENT” and “TENT w/ Align” are $\mathcal{L}_{ent}(\theta;\mathcal{X}^{t}_{i})$ and $\mathcal{L}_{ent}(\theta-\nabla\mathcal{L}_{con}(\theta;\mathcal{X}^{t}_{i});\mathcal{X}^{t}_{i})$, respectively.
Alignment between the empirical and pseudo gradients. To demonstrate that GraTa can align the empirical gradient and pseudo gradient more closely, we repeated the experiments on all scenarios and visualized the cosine similarity between these two gradients, where the empirical gradient is produced by the cross-entropy loss in a supervised manner, as shown in Figure 4. We reported the average results using each domain as the source domain and remaining ones as target domains. The greater the cosine similarity, the closer the pseudo gradient aligns with the empirical gradient. The results reveal the effectiveness of our GraTa in improving gradient direction and also indicate that even small alignment improvements can lead to significant performance gains.
@@
$\mathcal{L}_{ent}$ $\mathcal{L}_{con}$ $76.61$ $78.81$ $76.52$ $66.84$ $72.94$ 74.34
$\mathcal{L}_{rec}$ 76.76 $78.94$ $74.74$ $66.43$ $72.19$ $73.81$
$\mathcal{L}_{rot}$ $76.58$ 79.02 $74.39$ 67.90 $72.53$ $74.08$
$\mathcal{L}_{sup}$ $76.68$ $78.99$ $74.32$ $67.82$ $72.31$ $74.02$
$\mathcal{L}_{den}$ $76.71$ $78.94$ $75.11$ $65.18$ $72.26$ $73.64$
$\mathcal{L}_{con}$ $\mathcal{L}_{ent}$ $74.69$ $78.46$ $75.60$ $56.75$ $71.86$ $71.47$
$\mathcal{L}_{rec}$ $75.27$ $78.54$ $76.36$ $60.29$ $70.97$ $72.28$
$\mathcal{L}_{rot}$ $75.19$ $78.54$ $76.34$ $56.29$ $70.78$ $71.43$
$\mathcal{L}_{sup}$ $75.20$ $78.53$ $76.53$ $57.41$ $70.79$ $71.69$
$\mathcal{L}_{den}$ $75.37$ $78.52$ 76.67 $59.38$ $70.72$ $72.13$
$\mathcal{L}_{con}$ $\mathcal{L}_{rec}$ $75.18$ $77.85$ $75.13$ $62.33$ $73.71$ $72.84$
$\mathcal{L}_{ent}$ $74.47$ $77.74$ $74.71$ $59.64$ $73.49$ $72.01$
$\mathcal{L}_{rot}$ $75.74$ $78.04$ $75.12$ $60.13$ 73.91 $72.59$
$\mathcal{L}_{sup}$ $74.58$ $77.51$ $74.91$ $62.99$ $73.67$ $72.73$
$\mathcal{L}_{den}$ $75.44$ $77.82$ $75.66$ $60.12$ $73.73$ $72.55$
$\mathcal{L}_{con}$ $\mathcal{L}_{rot}$ $75.31$ $78.01$ $75.02$ $61.66$ $73.81$ $72.76$
$\mathcal{L}_{ent}$ $75.37$ $78.17$ $75.07$ $62.14$ $73.75$ $72.90$
$\mathcal{L}_{rec}$ $75.19$ $78.10$ $74.39$ $61.57$ $73.78$ $72.60$
$\mathcal{L}_{sup}$ $75.31$ $78.08$ $75.95$ $61.24$ $73.73$ $72.86$
$\mathcal{L}_{den}$ $75.27$ $78.21$ $74.84$ $61.56$ $73.65$ $72.70$
$\mathcal{L}_{con}$ $\mathcal{L}_{sup}$ $75.36$ $78.17$ $75.27$ $61.49$ $73.77$ $72.81$
$\mathcal{L}_{ent}$ $75.51$ $77.86$ $74.67$ $59.23$ $73.47$ $72.15$
$\mathcal{L}_{rec}$ $75.26$ $78.09$ $74.53$ $59.78$ $73.66$ $72.27$
$\mathcal{L}_{rot}$ $75.17$ $77.87$ $74.92$ $58.48$ $73.75$ $72.04$
$\mathcal{L}_{den}$ $74.94$ $77.91$ $74.83$ $60.65$ $73.63$ $72.39$
$\mathcal{L}_{con}$ $\mathcal{L}_{den}$ $75.34$ $78.02$ $74.47$ $61.74$ $73.71$ $72.66$
$\mathcal{L}_{ent}$ $75.20$ $77.44$ $74.77$ $59.22$ $73.38$ $72.00$
$\mathcal{L}_{rec}$ $75.40$ $78.09$ $74.51$ $60.71$ $73.59$ $72.46$
$\mathcal{L}_{rot}$ $75.32$ $77.71$ $76.49$ $55.30$ $73.44$ $71.65$
$\mathcal{L}_{sup}$ $75.48$ $77.87$ $75.08$ $59.82$ $73.68$ $72.38$
@@
More combinations of objective functions. For further discussion, we introduced four additional loss functions, i.e., the reconstruction loss, the rotation prediction loss, the super-resolution loss, and the denoising loss. We utilized the cross-entropy loss to calculate the rotation prediction loss and the mean-square-error loss to calculate others. To construct the super-resolution loss, we downsampled the input image by a factor of four. For the denoising loss, we added the Gaussian noise sampled from a standard normal distribution to the image. We evaluated all possible combinations of them to determine the optimal one and displayed the results in Table 5. It reveals that (1) using $\mathcal{L}_{con}$ as the pseudo objective function achieves superior performance over others; (2) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the pseudo objective function results in similar but suboptimal performance since they all require an additional branch to construct the loss, making it challenging provide appropriate gradients during inference; (3) using $\mathcal{L}_{rec}$, $\mathcal{L}_{rot}$, $\mathcal{L}_{sup}$, or $\mathcal{L}_{den}$ as the auxiliary objective function can enhance the performance of $\mathcal{L}_{con}$ and $\mathcal{L}_{ent}$, demonstrating the effectiveness of our GraTa.